# Conformalized Interactive Imitation Learning: Handling Expert Shift & Intermittent Feedback

### v1

@@ 5.1 Experimental Setup
Task & Initial Learner Policy.
The robot learns a neural network policy $\pi^{r}:\mathcal{X}\rightarrow\mathcal{A}$ to move a cup from a start to the expert’s desired goal location. We model input $x\in\mathcal{X}\subseteq\mathbb{R}^{12}$ as the xyz position of the robot as well as its binary gripper state (open or closed) across the previous 3 timesteps. The labels are actions, $a\in\mathcal{A}\subseteq\mathbb{R}^{4}$, represented as next xyz position and gripper state. The robot always starts at an initial $x$ with the gripper state closed (holding a cup) and must keep holding the cup while moving it to the unknown goal location that the expert prefers; let $\{g_{0},g_{1}\}$ be two such goals.
Before the first interactive deployment ($i=0$), we simulate the expert as initially giving demonstrations placing the cup at $g_{0}$, where $\pi^{h}_{i=0}(x)=x+\omega\frac{g_{0}-x}{\max_{d}(\mid g_{0}-x\mid)}$, where $\max_{d}(\cdot)$ is a maximum over dimensions. The step size is regulated to be at most $\omega=0.01$. The initial robot policy $\pi^{r}_{i=0}$ is trained on a dataset $\mathcal{D}_{0}$ of 10 expert trajectories with synthetically injected noise drawn from $\mathcal{N}(1,0.5)$ for robustness as in (Laskey et al., 2017). See Appendix D for policy implementation details.
Simulated Expert Policies. To induce controlled distribution shift, we study three expert policies (left, column of Figure 5). (1) Stationary: the expert has a fixed goal, $g_{0}$, across all deployment episodes. (2) Shift: the expert goal shifts from $g_{0}$ to $g_{1}$ at deployment episode $i=5$. For example, the expert may have decided that a different cup location is easier to reach.
(3) Drift: expert’s goal slowly drifts from $g_{0}$ to $g_{1}$ over the course of deployment episodes (in Figure 5 the drift from $g_{0}\rightarrow g_{1a}$ occurs at episode $i=5$, $g_{1a}\rightarrow g_{1b}$ occurs at episode $i=8$, and $g_{1b}\rightarrow g_{1}$ occurs at episode $i=11$). For example, the expert may start with a conservative goal location initially (e.g., a goal nearby) and incrementally move the cup closer and closer to their target goal that may be further out of reach.
@@
Methods. We compare ConformalDAgger to EnsembleDAgger (Menda et al., 2019).
ConformalDAgger uses an uncertainty threshold $\tau=0.09$, temperature $\beta=100$, lookback window $k=100$, $\text{lr}=0.6$, and initial $q^{lo,hi}_{0}=0.01$. The uncertainty threshold $\tau$ is heuristically tuned to ask few, but infrequent questions in the first interactive deployment episode. EnsembleDAgger queries an expert online when there is high action prediction variance across an ensemble of learner policies, and when a safety classifier detects dissimilarity between expert and robot actions. We use 3 ensemble members and an uncertainty threshold of $\tau=0.06$ for the ensemble disagreement (selected in a similar manner to ConformalDAgger) and a safety classifier threshold of $0.03$.
@@
Figure 5: Simulated Robot Results. When the expert shifts (middle row) or drifts (bottom), ConformalDAgger increases the number of requests for expert intervention compared to EnsembleDAgger, decreasing miscoverage and minimizing expert deviation after re-training. With a stationary expert (top row), both algorithms are similar. Shading represents std. error across 10 seeds.
@@ 5.2 Experimental Results
We focus results on infrequent human-gated feedback where $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=0.2,\forall t$. In Appendix D, we present experiments with frequent (=$0.9$) and partial (=$0.5$) feedback. Irrespective of the human-gated likelihood, the learner can always actively ask for feedback based on $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})$.
Takeaway 1: When the expert shifts, ConformalDAgger asks for more help immediately compared to EnsembleDAgger, enabling the algorithm to more quickly align to the expert. Consider deployment episode $i=5$ where the expert shifts from goal 0 to goal 1 (center row, Figure 5). ConformalDAgger immediately increases the number of expert feedback requests: before shift, the expert intervened $\sim$20of the time but in the 5th episode they intervene $\sim$60of the time. In contrast, EnsembleDAgger remains close to $\sim$30%. Relatedly, ConformalDAgger has consistently lower miscoverage rate (max = 0.4 at shift; converges to 0.2) compared to the baseline (max = 1.0 at shift, converges to 0.4). Due to the extra solicited feedback, ConformalDAgger’s ultimate decision and trajectory deviations are minimized in the subsequent retrained policies.
Takeaway 2: ConformalDAgger automatically asks for more help each time the expert drifts. In the bottom row of Figure 5, we see ConformalDAgger maintain a similar level of queries as EnsembleDAgger during the first shift (at episode 5) and last shift (episode 11) but increases feedback during the intermediate shift (at episode 8). Despite these similarities, EnsembleDAgger’s miscoverage rate is consistently higher than ConformalDAgger’s and EnsembleDAgger’s re-trained policy on average does not adapt as quickly (with higher expert decision and trajectory deviation). We hypothesize that this is because ConformalDAgger asks frequent questions consistently across all seeds compared to EnsembleDAgger (i.e. lower variance in Fig 5).
Takeaway 3: With a stationary expert, ConformalDAgger and EnsembleDAgger are similar. We find that without distribution shift (top row, Figure 5), both methods ask for minimal help (staying near the 20% human-gated probability), and achieve similar performance in alignment to the expert.
@@ 6 Hardware Experiments
Human Expert. The human expert teleoperates the robot via a Meta Quest 3 remote controller. They initially provide 100 demonstrated trajectories moving in a straight-line path along the marker line to construct the initial training dataset, $\mathcal{D}_{0}$. We model human-gated interventions with probability $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=0.2,~{}\forall t$. To control this rate, our interface queries the user with 20% probability at the any timepoint along deployment to simulate that consistent independent intervention rate.
We induce a potential distribution shift during interactive deployment episodes by changing the expert’s wiping strategy from a straight line to a zig-zag pattern.
@@
Learner Policy. We represent the robot’s policy as a CNN-based diffusion policy (Chi et al., 2023). The policy predicts $a^{r}\in\mathcal{A}\equiv\mathcal{Y}$ as 16 future actions, where each action is an end-effector position and quaternion orientation. The inputs $x\in\mathcal{X}$ are the current and previous image observations from the wrist and third-person camera. Image observations are encoded using a ResNet-18 visual encoder (trained end-to-end with the diffusion policy) and the action-generating process is conditioned on encoded observation features with FiLM (Perez et al., 2018). The initial policy $\pi^{r}_{0}$ is trained for 60K iterations (training parameters in Appendix Sec E.1).
The robot is deployed to interactively execute the task $50$ times after which its policy is retrained. This is an interactive deployment where the robot queries the user via our ConformalDAgger algorithm, or the user can independently intervene intermittently. During expert feedback, the human teleoperates the robot for a sequence of 16 timesteps, after which, if the robot doesn’t ask for feedback, control is handed back over to the robot.
After the interactive deployment, the initial learner policy is fine-tuned for an additional 60K iterations. We reset the learning-rate schedule, aggregate datapoints at which the expert gave feedback, and cap the training buffer size at 12k most recent input-action pairs.
ConformalDAgger Hyperparameters.
We apply IQT on only the predicted end-effector position, because we reason about nonconformity via the signed residual in Euclidean space. We initialize $q^{lo}_{0},q^{hi}_{0}=0.01$, our desired coverage level is $\alpha=0.1$, and $\gamma_{t}=0.15\hat{B}_{t}$, where the lookback window for $\hat{B}_{t}$ is 20.
@@
We use a simple model for $\mathbb{P}(\text{obs}^{r}\mid x_{t})=1$ that asks for help if the L2 norm of the uncertainty interval $C_{t}(x_{t})$ exceeds threshold $\tau=0.07$.
Figure 6: Hardware Robot Results. (left) ConformalDAgger increases the number of questions asked when the expert shifts in their strategy while Ensemble maintains a very similar question asking profile. (right) Qualitatively, ConformalDAgger can more effectively learn the new zigzag strategy.
Baseline.
Because the diffusion policy implicitly represents the action distribution present in the training data, we sample from the policy N=3 times and evaluate variance over the predicted action in order to capture the notion of epistemic uncertainty. We consider this an implicit EnsembleDAgger (Wolleb et al., 2022). Our implicit EnsembleDAgger implementation does not use the safety classifier.
EnsembleDAgger’s uncertainty threshold is 0.0005, tuned such that the learner asks occasional but not excessive questions. EnsembleDAgger and ConformalDAgger use the same initial learner policy.
Results.
@@
The left of Figure 6 shows quantitative results for both methods. ConformalDAgger increases the cumulative number of interventions up to about 100 timesteps of human feedback when interacting with the shifted expert (who switches to a zig-zag pattern) compared to the stationary expert (who always wipes in a straight line) which reaches a maximum of 40 timesteps. On the other hand, EnsembleDAgger maintains a very similar question-asking profile (Figure 6) for both the stationary and shifted expert, near 60 timesteps. To better understand this, we analyzed Figure 1, which shows ConformalDAgger engaging with each type of expert in one interactive deployment episode. Here, we see how ConformalDAgger does have uncertainty when interacting with the stationary expert (due to inherent noise in the VR-teleoperated intermittent human interventions) but the robot’s uncertainty never rises above the threshold where it initiates expert feedback. This is in stark contrast to the expert shift scenario (right of Figure 1) where half-way through the deployment episode the uncertainty intervals become large enough to trigger robot asking for help.
On the other hand, EnsembleDAgger requests more human interventions than ConformalDAgger at the start of the trajectory (as seen by the slope of cumulative interventions in Figure 6) but then plateaus.
We hypothesize that this may be because the training demonstrations have low variance while the sponge is in contact with the whiteboard, but higher variance as the robot approaches the table.
Because EnsembleDAgger’s uncertainty is uncalibrated to the deployment-time expert’s data, it is unable to identify the need for additional feedback when it reaches the whiteboard.
Qualitative results are shown on the right of Figure 6. As expected, with ConformalDAgger and EnsembleDAgger rollouts from the policies remain aligned with the stationary expert when retrained. However, under the shifted expert, rollouts from the retrained ConformalDAgger policy exhibit a more distinct zig-zag pattern, compared to EnsembleDAgger, likely because the EnsembleDAgger dataset has less datapoints because it does not ask for as much help.
@@ 7 Conclusion
In this work, we study uncertainty quantification for imitation learned policies that encounter distribution shift. We first extend uncertainty quantification via online conformal prediction to handle intermittent labels, such as those observed in interactive imitation learning. We then propose ConformalDAgger, a unification of our online conformal prediction algorithm with interactive imitation learning. Our approach provides asymptotic coverage guarantees for deployed end-to-end policies, uses the calibrated uncertainty measure to detect expert distribution shifts and actively query for more feedback, and empirically enables the robot learner update its policy to better align with the shifted expert distribution.
@@ Acknowledgments
The authors would like to thank Gokul Swamy for insightful conversations and the detailed review, Yilin Wu for help with diffusion policy and robot hardware setup, and Pat Callaghan for the detailed review. MZ is supported by an NDSEG fellowship.
@@
Appendix A Proof of Proposition 1

### v2

@@ 5.1 Experimental Setup
Task & Initial Learner Policy. The robot learns a neural network policy $\pi^{r}:\mathcal{X}\rightarrow\mathcal{A}$ to move a cup from a start to the expert’s desired goal location. We model input $x\in\mathcal{X}\subseteq\mathbb{R}^{12}$ as the xyz position of the robot as well as its binary gripper state (open or closed) across the previous 3 timesteps. The labels are actions, $a\in\mathcal{A}\subseteq\mathbb{R}^{4}$, represented as next xyz position and gripper state. The robot always starts at an initial $x$ with the gripper state closed (holding a cup) and must keep holding the cup while moving it to the unknown goal location that the expert prefers; let $\{g_{0},g_{1}\}$ be two such goals. Before the first interactive deployment ($i=0$), we simulate the expert as initially giving demonstrations placing the cup at $g_{0}$, where $\pi^{h}_{i=0}(x)=x+\omega\frac{g_{0}-x}{\max_{d}(\mid g_{0}-x\mid)}$, where $\max_{d}(\cdot)$ is a maximum over dimensions. The step size is regulated to be at most $\omega=0.01$. The initial robot policy $\pi^{r}_{i=0}$ is trained on a dataset $\mathcal{D}_{0}$ of 10 expert trajectories with synthetically injected noise drawn from $\mathcal{N}(1,0.5)$ for robustness as in (Laskey et al., 2017). See Appendix F for policy implementation details.
Simulated Expert Policies. To induce controlled distribution shift, we study three expert policies (left, column of Figure 5). (1) Stationary: the expert has a fixed goal, $g_{0}$, across all deployment episodes. (2) Shift: the expert goal shifts from $g_{0}$ to $g_{1}$ at deployment episode $i=5$. For example, the expert may have decided that a different cup location is easier to reach. (3) Drift: expert’s goal slowly drifts from $g_{0}$ to $g_{1}$ over the course of deployment episodes (in Figure 5 the drift from $g_{0}\rightarrow g_{1a}$ occurs at episode $i=5$, $g_{1a}\rightarrow g_{1b}$ occurs at episode $i=8$, and $g_{1b}\rightarrow g_{1}$ occurs at episode $i=11$). For example, the expert may start with a conservative goal location initially (e.g., a goal nearby) and incrementally move the cup closer and closer to their target goal that may be further out of reach. (4) Environment Shift: Under environment shift, the policy experiences covariate shift, where the deployment-time state distribution differs from the training distribution. Because the policy is trained on inputs which capture the robot’s position, we instantiated environment shift by changing the starting position of the robot in the first episode, maintaining this new position for all latter episodes.
@@
Methods. We compare ConformalDAgger to EnsembleDAgger (Menda et al., 2019), LazyDAgger (Hoque et al., 2021b), and SafeDAgger (Zhang & Cho, 2017). We performed a hyperparameter sensitivity analysis for all methods (see B) and selected thresholds such that all approaches ask infrequent questions in the first few interactive deployment episodes, where no shift has occurred. ConformalDAgger uses an uncertainty threshold $\tau=0.06$, temperature $\beta=100$, lookback window $k=100$, $\text{lr}=0.6$, and initial $q^{lo,hi}_{0}=0.01$. The uncertainty threshold $\tau$ is heuristically tuned to ask few, but infrequent questions in the first interactive deployment episode. EnsembleDAgger queries an expert online when there is high action prediction variance across an ensemble of learner policies, and when a safety classifier detects dissimilarity between expert and robot actions. We use 3 ensemble members and an uncertainty threshold of $\tau=0.06$ for the ensemble disagreement and a safety classifier threshold of $s=0.03$. LazyDAgger uses $s=0.03$ to begin human intervention, and only switches back to autonomous mode when the deviation between the learner’s prediction and expert’s action are below a context-switching threshold, $0.1*s$. To make SafeDAgger’s number of initial interventions comparable, we decrease the safety classifier threshold to $s=0.01$.
@@
Figure 5: Simulated Robot Results. When the expert shifts or drifts (middle rows), ConformalDAgger increases the number of requests for expert feedback compared to EnsembleDAgger and SafeDAgger, but has less interventions than LazyDAgger. This decreases miscoverage and expert deviation after re-training. With a stationary expert (top) or environment shift (bottom), both EnsembleDAgger and ConformalDAgger perform similarly. Shading is std. error across 5 seeds.
@@ 5.2 Experimental Results
We focus results on infrequent human-gated feedback where $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=0.2,\forall t$. In Appendix F, we present experiments with frequent (=$0.9$) and partial (=$0.5$) feedback. Irrespective of the human-gated likelihood, the learner can always actively ask for feedback based on $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})$.
Takeaway 1: When the expert shifts, ConformalDAgger asks for more help immediately compared to EnsembleDAgger and SafeDAgger, enabling the algorithm to more quickly align to the expert with fewer human interventions than LazyDAgger. Consider deployment episode $i=5$ where the expert shifts from goal 0 to goal 1 (center row, Figure 5). ConformalDAgger immediately increases the number of expert feedback requests: before shift, the expert intervened $\sim$20of the time but in the 5th episode they intervene $\sim$60of the time. In contrast, EnsembleDAgger remains close to $\sim$30%. Relatedly, ConformalDAgger has consistently lower miscoverage rate (max = 0.4 at shift; converges to 0.2) compared to the baseline (max = 1.0 at shift, converges to 0.4). Due to the extra solicited feedback, ConformalDAgger’s ultimate decision and trajectory deviations are minimized in the subsequent retrained policies.
Takeaway 2: ConformalDAgger automatically asks for more help each time the expert drifts. In the third row of Figure 5, we see ConformalDAgger maintain a similar level of queries as EnsembleDAgger during the first shift (at episode 5) and last shift (episode 11) but increases feedback during the intermediate shift (at episode 8). Despite these similarities, EnsembleDAgger’s miscoverage rate is consistently higher than ConformalDAgger’s and EnsembleDAgger’s re-trained policy on average does not adapt as quickly (with higher expert decision and trajectory deviation). We hypothesize that this is because ConformalDAgger asks frequent questions consistently across all seeds compared to EnsembleDAgger (i.e. lower variance in Fig 5).
Takeaway 3: With a stationary expert, ConformalDAgger and baselines are similar. We find that without distribution shift (top row, Figure 5), all methods ask for minimal help (staying near the 20% human-gated probability), with LazyDAgger asking slightly more questions, and achieve similar performance in alignment to the expert.
Takeaway 4: Under environment shift, ConformalDAgger is comparable to baselines. In the fourth row of Figure 5, we see that EnsembleDAgger and ConformalDAgger balance intervention frequency with policy learning quality marginally better than LazyDAgger or SafeDAgger.
@@ 6 Hardware Experiments
Human Expert. The human expert teleoperates the robot via a Meta Quest 3 remote controller. They initially provide 50 demonstrated trajectories moving in a straight-line path along the marker line to construct the initial training dataset, $\mathcal{D}_{0}$. We model human-gated interventions with probability $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=0.2,~{}\forall t$. To control this rate, our interface queries the user with 20% probability at the any timepoint along deployment to simulate that consistent independent intervention rate. We induce a potential distribution shift during interactive deployment episodes by changing the expert’s wiping strategy from a straight line to a zig-zag pattern.
@@
Learner Policy. We represent the robot’s policy as a CNN-based diffusion policy (Chi et al., 2023). The policy predicts $a^{r}\in\mathcal{A}\equiv\mathcal{Y}$ as 16 future actions, where each action is an end-effector position and quaternion orientation. The inputs $x\in\mathcal{X}$ are the current and previous image observations from the wrist and third-person camera. Image observations are encoded using a ResNet-18 visual encoder (trained end-to-end with the diffusion policy) and the action-generating process is conditioned on encoded observation features with FiLM (Perez et al., 2018). The initial policy $\pi^{r}_{0}$ is trained for 50K iterations (training parameters in Appendix Sec H.1). The robot is deployed to interactively execute the task $25$ times after which its policy is retrained. This is an interactive deployment where the robot queries the user via our ConformalDAgger algorithm, or the user can independently intervene intermittently. During expert feedback, the human teleoperates the robot for a sequence of 16 timesteps, after which, if the robot doesn’t ask for feedback, control is handed back over to the robot. After the interactive deployment, the initial learner policy is fine-tuned for an additional 35K iterations. We reset the learning-rate schedule, aggregate datapoints at which the expert gave feedback, and cap the training buffer size at 6k most recent input-action pairs.
ConformalDAgger Hyperparameters. We apply IQT on only the predicted end-effector position, because we reason about nonconformity via the signed residual in Euclidean space. For quaternion rotation, we calibrate only the positional output to avoid the complexity of multiple IQT processes for positional and rotational nonconformities.
@@
We initialize $q^{lo}_{0},q^{hi}_{0}=0.01$, our desired coverage level is $\alpha=0.1$, and $\gamma_{t}=0.15\hat{B}_{t}$, where the lookback window for $\hat{B}_{t}$ is 20. We use a simple model for $\mathbb{P}(\text{obs}^{r}\mid x_{t})=1$ that asks for help if the L2 norm of the uncertainty interval $C_{t}(x_{t})$ exceeds threshold $\tau=0.07$.
Figure 6: Hardware Results. (bottom, left) Conformal enables the robot to significantly increase the number of questions when the expert shifts their strategy while Safe and Ensemble only increase slightly. (bottom, right) Qualitatively, Conformal learns a higher quality zigzag strategy. Lazy exhibits a similar an intervention profile to Conformal, but because the algorithm doesn’t ask for help consistently in each episode, it does not receive enough additional data to learn the shifted policy.
Baseline. Because the diffusion policy implicitly represents the action distribution present in the training data, we sample from the policy N=3 times and evaluate variance over the predicted action in order to capture the notion of epistemic uncertainty. We consider this an implicit EnsembleDAgger: we use the implicit ensemble approach from (Wolleb et al., 2022) by sampling multiple times from the same diffusion model. The safety classifier is a fully connected network that takes as input the ResNet-18 embeddings of the camera images concatenated with robot proprioceptive state. We set the classifier threshold $s$ to be the mean of the distribution of expert-learner divergence over the training datapoints. EnsembleDAgger’s uncertainty threshold is 0.005, at approximately the mean of the distribution of ensembles variances (over position and orientation predicted by the policy) over the training data, tuned such that the learner asks occasional but not excessive questions. We also compared with LazyDAgger (threshold of $0.25*s$, informed by Hoque et al. (2021a)), and SafeDAgger. All models use the same initial learner policy.
@@
Results. The left of Figure 6 shows quantitative results for both methods. ConformalDAgger increases the cumulative number of robot-requested interventions up to about 35 timesteps of human feedback when interacting with the shifted expert (who switches to a zig-zag pattern) compared to the stationary expert (who always wipes in a straight line) which does not induce as frequent of questions. To better understand this, we analyzed Figure 1, which shows ConformalDAgger engaging with each type of expert in one interactive deployment episode. Here, we see how ConformalDAgger has uncertainty when interacting with the stationary expert (due to inherent noise in the VR-teleoperated intermittent human interventions) but the robot’s uncertainty never rises above the threshold the robot initiates expert feedback. This is in stark contrast to the expert shift scenario (right of Figure 1) where half-way through the deployment episode the uncertainty intervals become large enough to trigger robot asking for help. On the other hand, with the shifted expert, EnsembleDAgger and SafeDAgger do not request as many interventions as LazyDAgger and ConformalDAgger, (Figure 6). When LazyDAgger asks for help the queries follow a similar profile as ConformalDAgger, but it does not reliably ask every episode, leading to less data for learning the shifted strategy. We hypothesize that this may be because the training demonstrations have low variance while the sponge is in contact with the whiteboard, but higher variance as the robot approaches the table. Because EnsembleDAgger’s and SafeDAgger’s uncertainty is uncalibrated to the deployment-time expert’s data, it is unable to identify the need for additional feedback when it reaches the whiteboard. Qualitative results are shown on the right of Figure 6. As expected, all approaches remain aligned with the stationary expert when retrained. However, under the shifted expert, rollouts from the retrained ConformalDAgger policy exhibit a more distinct zig-zag pattern, compared to EnsembleDAgger and SafeDAgger, likely because the approaches did not ask for as much help.
@@ 7 Conclusion
We first extend uncertainty quantification via online conformal prediction to handle intermittent labels, such as those observed in interactive imitation learning. We then propose ConformalDAgger, a unification of our online conformal prediction algorithm with interactive imitation learning. Our approach provides asymptotic coverage guarantees for deployed end-to-end policies, uses the calibrated uncertainty measure to detect expert distribution shifts and actively query for more feedback, and empirically enables the robot learner update its policy to better align with the expert.
@@ Acknowledgments
The authors would like to thank Gokul Swamy for insightful conversations and the detailed review, Yilin Wu for help with diffusion policy and robot hardware setup. MZ is supported by an NDSEG fellowship.
@@
Appendix A Additional Background on Conformal Predicition for Robotics
Appendix B Hyperparameter Sensitivity Analysis
Figure 7: Hyperparameter sensitivity analysis under Expert Shift (Human-gated probability $P(obs^{h}_{t}|o_{t})=0.2$).
We show below a sensitivity analysis in the Expert Shift (Human-gated probability $P(obs^{h}_{t}|o_{t})=0.2$) context of the uncertainty threshold for ConformalDAgger and EnsembleDAgger, where we vary the uncertainty threshold between 0.03 and 0.09 (Figure 7). To demonstrate the results over the deployment episodes, we hold the EnsembleDAgger safety threshold constant at 0.03. The results are averaged over 3 seeds.
When the threshold is low (0.03), the approaches all identify the shift and ask for help accordingly; however, this comes at the cost of a great deal of human feedback requested throughout every deployment episode. Even in the first episode in which there has been no expert shift, a low-threshold at 0.03 requires conformal and ensemble to ask questions 70-80% of the time. When the threshold is set too high, at 0.09, EnsembleDAgger is unable to identify the shift and continuously mispredict.
We included in our sensitivity analysis LazyDAgger ’s context-switching threshold, where we tested values of 0.1, 1, and 10, which are thresholds presented in the LazyDAgger sensitivity analysis in Hoque et al. (2021b). Similarly, we hold the LazyDAgger safety threshold constant at 0.03. For SafeDAgger, we examined safety thresholds of 0.01, 0.03, and 0.05.
Across all of the algorithms, ConformalDAgger’s expert deviation is the least prone to threshold selection, where all models are able to effectively adapt to the expert’s shifted policy. Empirically, we find that if $\tau$ is set too high (at 0.18, 0.21), ConformalDAgger may not flag high uncertainty, relying instead on human-gated interventions to update IQT. When we ablate the values of tau between the values [0.03, 0.06, 0.09, 0.12, 0.15], we found ConformalDAgger to be fairly robust in terms of miscoverage (see updated Figure 7 in Appendix B, $\tau$=0.18, 0.21 included as well).
Theoretically, if $\tau$ is set too high, failing to properly reflects the distribution of encountered misprediction residuals, ConformalDAgger will not flag high uncertainty, relying on human-gated interventions only ($P(obs^{h}_{t}|o_{t})$) to update IQT and provide retraining data, which may result in poor coverage. However, importantly, one benefit of IQT is that its theoretical guarantees afford us an understanding of what the miscoverage gap will be. Recall that Proposition 1 gives us a bound on the miscoverage gap between the achieved miscoverage rate and desired miscoverage rate $\alpha$. This gap depends on 1) the time horizon $T$ that we are running our algorithm, and 2) the maximum value of $\frac{\gamma_{t}}{p_{t}}$. In cases where $\tau$ is set too high such that ($P(obs^{r}_{t}|o_{t})$ is close to 0), meaning the robot almost never queries for additional help, the probability of getting feedback $p_{t}$ is approximately the $P(obs^{h}_{t}|o_{t})$, relying only on the likelihood of human feedback. For short sequences ($T$ small), and infrequent human-gated feedback ($p_{t}$ small), Proposition 1 informs us that this miscoverage gap might be large. Proposition 1 shows the theoretical basis for that when human-gated feedback is infrequent, and highlights that a proper and effective selection of $\tau$ to raise $P(obs^{r}_{t}|o_{t})$ and subsequently $p_{t}$, helps to reduce the miscoverage gap.
The red arrows indicate the threshold values we ended up using for our 4D reaching experiments. We used the number of human interventions requested as a proxy metric for helping us choose the thresholds. Although the EnsembleDAgger threshold of 0.03 more quickly learns the expert policy than 0.06, however, the 0.03 threshold causes the robot to asks many more questions at the initial episode (even without the existence of shift). We chose the threshold of 0.06 for EnsembleDAgger and ConformalDAgger because in the first 4 episodes before the shift occurs, the algorithm asks some, but not excessive, questions. LazyDAgger uses the safety classifier prediction, $s=0.03$ to begin human intervention, and only switches back to autonomous mode when the learner’s prediction and human ground truth action received online are below a context-switching threshold, $0.1s$. Because SafeDAgger alone under $s=0.03$ does not ask as many questions as EnsembleDAgger and LazyDAgger which both involve additional mechanisms for querying alongside the safety classifier, we choose $s=0.01$ for SafeDAgger.
Appendix C Proof of Proposition 1